A rocket moves straight upward, starting from rest with an acceleration of 29.4 m/s2. It runs out of fuel at the end of 4s and continues to coast upward, reaching a max height before falling back to Earth. 

(a) Find the rocket’s velocity and 

position at the end of 4 s. 

(b) Find the rocket’s max height.

(c) Find it’s velocity before crashing to the ground.



1
Expert's answer
2020-10-05T10:52:40-0400

a) We can find the rocket's velocity and position from the kinematic euations:


"v = v_0 + at,""y = v_0t + \dfrac{1}{2}at^2,"

here, "v_0 = 0 \ \dfrac{m}{s}" is the initial velocity of the rocket, "a = 29.4 \ \dfrac{m}{s^2}" is the acceleration of the rocket, "t = 4 \ s" is the time during which the rocket runs out of fuel.

Then, we get:


"v(t=4 \ s) = 29.4 \ \dfrac{m}{s^2} \cdot 4 \ s = 117.6 \ \dfrac{m}{s},""y(t=4 \ s) = \dfrac{1}{2} \cdot 29.4 \ \dfrac{m}{s^2} \cdot (4 \ s)^2 = 235.2 m."

b) Let's first find the distance traveled by the rocket from the end of 4s until it reaches max height:


"v_f^2 = v_i^2 + 2gy_2,"

here, "v_f = 0 \ \dfrac{m}{s}" is the final velocity of the rocket, "v_i = 117.6 \ \dfrac{m}{s}" is the initial velocity of the rocket, "g = -9.8 \ \dfrac{m}{s^2}" is the acceleration due to gravity, "y_2" is the distance traveled by the rocket from the end of 4s until it reaches max height.

Then, from this equation we can find "y_2":


"y_2= \dfrac{v_f^2-v_i^2}{2a}=\dfrac{(0 \ \dfrac{m}{s})^2-(117.6 \ \dfrac{m}{s})^2}{2 \cdot (-9.8 \ \dfrac{m}{s^2})} = 705.6 \ m."

Finally, we can find the rocket's max height:


"h_{max} = y_1 + y_2,"

here, "y_1 = 235.2 \ m" is the distance traveled by the rocket at the end of 4s.

Therefore, we get:


"h_{max}= 235.2 \ m + 705.6 \ m = 941 \ m."

c) We can find the rocket velocity before crashing to the ground from the kinematic equation:


"v_f^2 = v_i^2 + 2gh_{max},""v_f = \sqrt{v_i^2 + 2gh_{max}},""v_f = \sqrt{(0 \ \dfrac{m}{s})^2 + 2 \cdot 9.8 \ \dfrac{m}{s^2} \cdot 941 \ m} = 135.8 \ \dfrac{m}{s}."

Answer:

a) "v(t=4 \ s) = 117.6 \ \dfrac{m}{s}, y(t=4 \ s) = 235.2 m."

b) "h_{max}=941 \ m."

c) "v_f =135.8 \ \dfrac{m}{s}."